OpenStudy (anonymous):

the demand for a product is given by p=108−0.015x dollars, where p is the price per unit and x is the number of units. The cost of producing x units is given by C=660+90x dollars, and the profit for producing x units is given by P=R−C=xp−C. Skech the graph of the profit function and estimate the number of units that would produce a maximum profit. x for maximum

OpenStudy (jdoe0001):

well, P = R- C or P = x(p) -C so... get x(p) -C and see what you get for the equation

OpenStudy (anonymous):

-0.015x^2+18x-660

OpenStudy (anonymous):

i got the answer! Thanks :) it was 600